# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v1

@@ Discussion
Figure 5: Simulation fringe pattern fit. This figure compares the simulation results of both methods, the top for PSI and the bottom for SPSI. The light blue shaded area depicts the total number of atoms in each bin, denoted as $N_{\text{total}}$, while the darker shade indicates only those atoms in the internal state $|1\rangle$, denoted as $N_{|1\rangle}$. The green points represent the normalized probability for the $|1\rangle$ state, calculated as $P_{1}=N_{|1\rangle}(x)/N_{\text{total}}(x)$, while the red line denotes the sinusoidal fit to these green points. The angular velocity is determined from the fit’s fringe spatial frequency using Eq. 3. The simulations begin with an initial 1D Gaussian cloud comprising $N=10^{6}$ atoms, with a width of $\sigma_{x0}=100\,\rm{\mu m}$ and a temperature of $T=5\,\rm{\mu K}$. For the PSI method, the atoms undergo free expansion during the interferometer sequence, with a time interval between pulses of $T_{R}=25\,\rm ms$, while the simulated angular velocity is $\Omega=40\,\rm{mrad/s}$. In the case of SPSI, the cloud is additionally subjected to a repulsive potential having a power of $P=1\,\rm{W}$, with a beam cross-section of $400\crossproduct 400\,\mu$m2, activated for $t_{\text{acc}}=20\,\rm{\mu s}$. This results in a potential frequency of $\omega=2\pi\times 1.214\,\rm{kHz}$ (Eq. 11), a squeezing parameter $\eta=5.4$ (Eq. 8), and an effective time $t_{\text{eff}}=0.84\,\rm{ms}$ (Eq. 9). Consequently, the cloud size increases significantly. Note that the contrast and number of fringes in the PSI method is smaller than that of the SPSI (see Figs. 1-2), leading to a fit deviation and uncertainty one order of magnitude higher for the PSI case.
@@
The normalized population is fitted to a sinusoidal function, $f(x)=A\sin(Bx+C)+D$, which initially employs an FFT algorithm to identify optimal frequencies and uses them as initial guesses for the fit algorithm, based on non-linear least-squares. To improve the fit’s success rate, we implement a basic image processing stage before normalizing $P_{1}$. This stage involves applying different averaging windows to the image’s pixels to reduce noise in the atom count in each bin. Additionally, the algorithm examines various windows of interest from the image, excluding the noisier edges where fewer atoms are counted and more significant fluctuations occur.

### v2

@@ Discussion
Figure 5: Simulation fringe pattern fit. This figure compares the simulation results of both methods, the top for PSI and the bottom for SPSI. The light blue shaded area depicts the total number of atoms in each bin, denoted as $N_{\text{total}}$, while the darker shade indicates only those atoms in the internal state $|1\rangle$, denoted as $N_{|1\rangle}$. The green points represent the normalized probability for the $|1\rangle$ state, calculated as $P_{1}=N_{|1\rangle}(x)/N_{\text{total}}(x)$, while the red line denotes the sinusoidal fit to these green points. The angular velocity is determined from the fit’s fringe spatial frequency using Eq. 3. The simulations begin with an initial 1D Gaussian cloud comprising $N=10^{6}$ atoms, with a width of $\sigma_{x0}=100\,\rm{\mu m}$ and a temperature of $T=5\,\rm{\mu K}$. For the PSI method, the atoms undergo free expansion during the interferometer sequence, with a time interval between pulses of $T_{R}=25\,\rm ms$, while the simulated angular velocity is $\Omega=40\,\rm{mrad/s}$. In the case of SPSI, the cloud is additionally subjected to a repulsive potential having a power of $P=1\,\rm{W}$, with a beam cross-section of $400\crossproduct 400\,\mu$m2, activated for $t_{\text{acc}}=20\,\rm{\mu s}$. This results in a potential frequency of $\omega=2\pi\times 1.214\,\rm{kHz}$ (Eq. 11), a squeezing parameter $\eta=5.4$ (Eq. 8), and an effective time $t_{\text{eff}}=0.84\,\rm{ms}$ (Eq. 9). Consequently, the cloud size increases notably. Note that the contrast and number of fringes in the PSI method is smaller than that of the SPSI (see Figs. 1-2), leading to a fit deviation and uncertainty one order of magnitude higher for the PSI case.
@@
The normalized population is fitted to a sinusoidal function, $f(x)=A\sin(Bx+C)+D$, which initially employs an FFT algorithm to identify optimal frequencies and uses them as initial guesses for the fit algorithm, based on non-linear least-squares. To improve the fit’s success rate, we implement a basic image processing stage before normalizing $P_{1}$. This stage involves applying different averaging windows to the image’s pixels to reduce noise in the atom count in each bin. Additionally, the algorithm examines various windows of interest from the image, excluding the noisier edges where fewer atoms are counted and more fluctuations occur.